OpenStudy (anonymous):

the lengths of the corresponding sides of two similar trapezoids are in the ratio 3:5. if the area of the smaller trapezoid is 63 square inches, what is the area of the larger trapezoid?

OpenStudy (anonymous):

Hint: area is proportional to the square of perimeter for any shape.

OpenStudy (anonymous):

I need a bit more

OpenStudy (anonymous):

If the ratio of corresponding sides is \(3:5\), then what is the ratio of the perimeters?

OpenStudy (anonymous):

You've lost me. I don't understand this chapter at all

OpenStudy (anonymous):

Try looking at it this way: When you double the perimeter of a shape by scaling it, you will quadruple the area. When you triple the perimeter of a shape by scaling it, you will have nine times the area. When you scale it to four times the perimeter, you will have 16 times the area. Keeping this pattern in mind, you can reason out how many times bigger or smaller the area of the new trapezoid will be relative to the old one.

Directrix (directrix):

@jfron18 I will attach the theorem you need to learn to be able to work this type problem. (3/5) ^ 2 = 63/A where A is the area of the larger trapezoid 9/25 = 63/A 9A = 63 * 25 A = ( 63 * 25) / 9 A = ?

Directrix (directrix):

If two similar triangles have a scale factor of a : b, then the ratio of their areas is a^2: b^2.